OpenStudy (surryyy):

Determine f(5) for f(x) = 1/3 (5 - 4x). 5 -55/3 25/3 -5

OpenStudy (cggurumanjunath):

put x = 5

OpenStudy (anonymous):

yeah just substitute x as 5 and solve

OpenStudy (anonymous):

do you need help solving it still?

OpenStudy (surryyy):

it would be -5?

OpenStudy (anonymous):

yup Cx

OpenStudy (anonymous):

1/3[5-4(5)] 1/3[5-20] 1/3[-15] -5

OpenStudy (surryyy):

lol thnx.

OpenStudy (jhannybean):

\[\frac{1}{3} (5-4(\color{red}{5}))\]\[\frac{1}{3} (5-20) = \frac{1}{3}(-15)= -5\]